OpenStudy (nuccioreggie):

Which of these is a simplified form of the equation 6y + 4 = 8 + 2y + 2y? 6y = 8 10y = 12 2y = 4 4y = 12

OpenStudy (tkhunny):

Have you considered rearranging by addition and subtraction?

OpenStudy (nuccioreggie):

yes

OpenStudy (whovianchick):

6y+4=8+2y+2y 6y+4 = 8 +4y (combine like terms) 2y+4 =8 (subtract 4y from both sides) 2y= 4 (subtract 4 from both sides) C: 2y=4

OpenStudy (mathmale):

"Rearranging," for the purpose of consolidating (grouping together) "like" terms: group together all "y" terms and all constant terms.

OpenStudy (whovianchick):

The answer is C: 2y=4